### v1

# Towards Intelligent Battery Management via A Five-Tier Digital Twin Framework

Abstract
Battery management systems (BMSs) are critical to ensuring safety, efficiency, and longevity across electronics, transportation, and energy storage. However, with the rapid growth of lithium-ion batteries, conventional reactive BMS approaches face limitations in health prediction and advanced maintenance management, resulting in increased safety risks and economic costs. To address these challenges, we propose a five-tier digital twin framework for intelligent battery management. The framework spans geometric visualization, predictive modeling, prescriptive optimization, and autonomous operation, enabling full lifecycle optimization. In validation, an electrochemical model calibrated via Bayesian optimization achieved strong alignment with measured voltage and temperature, with Mean Absolute Percentage Errors (MAPE) below 1.57% and 0.39%. A Physics-Informed Neural Network (PINN) then combined data and simulations to predict State of Health (SOH), attaining MAPE under 3% with quantified uncertainty. This framework elevates BMSs into intelligent systems capable of proactive management and autonomous optimization, advancing safety and reliability in critical applications.
Index Terms:
Digital twin, Battery Management System (BMS), Bayesian optimization, Physics-informed Neural Network (PINN), State of Health (SOH) prediction
I Introduction
Battery management systems (BMSs) are critical for ensuring the safety, efficiency, and longevity of batteries in Electric Vehicles (EVs) and grid energy storage. The indispensable role of BMSs is driven by the rapid expansion of energy storage application, expected to reach around 442 GWh globally by 2030 [1]. This accelerating deployment is not only increasing capacity demands but also introducing growing system complexity. For instance, EV battery packs contain thousands of cells operating under fluctuating loads, while grid-scale systems integrate heterogeneous batteries with diverse degradation patterns. These complexities demand BMSs capable of coordinating cell behavior and maintaining resilience under varying demand and supply conditions.
Conventional BMSs face significant challenges in managing dynamic operational conditions, with limited predictive capabilities and inadequate sensing that restricts the modeling of critical internal states. They typically rely on easily measurable signals such as current, terminal voltages, and surface temperatures, which provide limited visibility into latent states like lithium concentration gradients, internal temperature distributions, State of Charge (SOC), or State of Health (SOH) [2]. Without such detailed information, it becomes difficult to link sensor data to underlying degradation mechanisms or to optimize charge–discharge strategies.
Most existing BMS implementations remain rule-based designs, in which threshold-triggered protective actions provide essential short-term safety. This design arises because general BMS architectures rely on embedded Programmable Logic Controllers (PLCs) and a central Electronic Control Unit (ECU), which process only a limited set of measurable signals [3]. These controllers can activate local interlocks, shutdowns, or alarms to ensure immediate responses against incidents like short circuits or thermal runaway. While effective for mitigating immediate operational risks, this inherently reactive approach offers minimal predictive insights into long-term battery health and degradation mechanisms. To enable a more active, adaptive, and intelligent BMS, recent research has explored the integration of battery digital twins. A digital twin is a synchronized digital replica of the physical system that combines the asset itself, a high-fidelity virtual model, and bi-directional data flows to maintain alignment [4]. In batteries, digital twins often leverage electrochemical and thermal modeling, supported by remote monitoring and cloud-based computation, to deliver improved insights into internal states and predictive control [5].
Despite existing advancements in both traditional BMS and current digital twin solutions, there are several technical gaps mentioned above are still remain and need to be overcomed. First, existing research lacks a digital twin framework that tightly integrates high-fidelity multi-physics battery models with advanced AI techniques. Current battery digital twin approaches either rely on simplified physics or purely data-driven surrogates, which compromises both accuracy and generalization, making the virtual replica less interpretable and trustworthy. Second, the absence of a unified system-level framework means that the perception–prediction–control loop is handled in isolation, leading to fragmented insights and delayed responses. Consequently, current digital twin systems struggle to keep operations and virtual models aligned with the accuracy and speed needed for autonomous decision-making.
To bridge these gaps, we propose a five-tier digital twin framework systematically evolving battery management from geometric modeling to autonomous control. The framework integrates real-time data assimilation at the descriptive tier, multi-physics simulation for accurate forecasting at the predictive tier, optimization algorithms for prescriptive control, and autonomous closed-loop operation at the highest tier, offering unprecedented predictive accuracy and operational intelligence. To realize these functions, the framework is validated through a co-simulation methodology. Electrochemical and thermal dynamics are modeled in physics-based simulators and synchronized within NVIDIA Omniverse, which provides a unified 3D virtual environment. Continuous-time prediction is supported by NVIDIA PhysicsNeMo to train physics-informed models that rapidly simulate real-world battery behaviors across diverse operating conditions. Reinforcement learning–based adaptation supports optimization under dynamic conditions, and closed-loop feasibility is tested through Hardware-in-the-Loop (HIL) validation. Different from conventional designs that treat monitoring, forecasting, and actuation as separate modules, this architecture establishes a unified digital intelligence pipeline evolving across the entire battery lifecycle, enabling interpretable modeling, reliable prediction, and intelligent control.
The contributions of our approach are as follows.
  • To the best of our knowledge, this is the first article that systematically proposes and introduces a unified five-tier digital twin framework for intelligent battery management. Each tier is clearly defined to address the fundamental gaps in existing BMS approaches.
  • To the best of our knowledge, this is the first framework that employs calibrated multi-physics modeling with PIML to achieve interpretable tracking of internal battery states and latent variables, which directly enhances the physical interpretability of battery management and supports proactive decision-making.
  • We further validate the effectiveness of our framework through a comprehensive case study on multi-physics model calibration and SOH prediction. Our approach achieves remarkable accuracy, reducing voltage and temperature simulation Mean Absolute Percentage Error (MAPE) to less than 1% through Bayesian optimization. Moreover, our physics-informed neural networks combined with uncertainty quantification methods achieve robust SOH predictions with a MAPE consistently below 3%, demonstrating substantial improvements over conventional battery management methods.
II PRELIMINARIES
In this section, we first review BMS architectures and the limitations of conventional rule-based designs. We then introduce multi-physics modeling approaches for electrochemical, thermal, and degradation processes. Finally, we summarize key simulation platforms and validation methods that bridge virtual models with real-world operation.
II-A Overview of BMS
BMSs are designed to monitor, protect, and optimize batteries. In practice, a BMS integrates subsystems for signal acquisition, central control, circuit protection, and communication with upper-level energy management systems [6]. Despite these capabilities, conventional implementations remain limited in intelligence, as they rely heavily on rule-based logic where protective actions such as disconnection, cooling, or cell balancing are triggered only after predefined safety thresholds are exceeded [7]. hile such designs mitigate immediate risks, they remain inherently reactive and inefficient, motivating the need for more advanced and predictive BMS solutions.
II-B Battery Multi-physics Modeling Techniques
Battery multi-physics modeling sets up the foundation for estimating and predicting battery dynamic behavior based on electrochemical, chemical and mechanical principles. At the electrochemical level, broadly used models include the Single Particle Model (SPM) and the Doyle–Fuller–Newman (DFN) model are developed to model the temperature, cell voltage and other critical states [8]. The SPM simplifies battery dynamics by considering single representative particles, offering computational efficiency, while the DFN model provides detailed insights by simulating complex lithium-ion transport and electrochemical reactions across battery electrodes.
Thermal modeling is a key component that characterizes heat generation, transfer, and dissipation during battery operation, typically formulated from energy conservation laws. Lumped thermal models [9] are widely adopted due to their simplicity and computational efficiency in estimating temperature dynamics, thereby supporting effective thermal management strategies essential for safety and reliability. Battery degradation modeling, on the other hand, focuses on long-term aging driven by mechanisms such as Solid Electrolyte Interphase (SEI) growth, lithium plating, and particle cracking. Advanced frameworks, such as those introduced by Wang et al. [10], explicitly couple these processes to capture the complex interactions that govern performance decay and material loss. These integrated multi-physics modeling plays an important role in accurately predicting battery lifetime and performance, emphasizing the importance and challenge of incorporating comprehensive multi-physics models within digital twin architectures. It is worth mentioning that both the thermal models and degradation models can be coupled within the multi-physics model, such as SPM and DFN, by introducing new source functions.
II-C Simulation Platforms and Validation Methods
High-fidelity simulation platforms are a trustworthy way to validate battery digital twins, ensuring accurate representation of electrochemical, thermal, and mechanical behaviors under diverse conditions. Besides well-known tools like PyBaMM and COMSOL, other notable platforms include ANSYS Fluent for computational fluid dynamics and thermal analysis, and MATLAB/Simulink for dynamic system modeling and control strategy simulations. These platforms offer great support in analyzing and simulating battery operation, facilitating comprehensive case study evaluation and virtual testing without the necessity for costly physical prototype validations.
HIL and Software-in-the-Loop (SIL) are two main complementary simulation paradigms that establish a bridge between virtual models and real-world systems [11]. In SIL, controller algorithms and software modules are embedded within a virtual simulation environment, allowing rapid prototyping, algorithm debugging, and iterative design without the need for physical hardware. Moreover, HIL extends this principle by incorporating real physical components into the simulation loop, thereby exposing algorithms to realistic operating conditions and hardware constraints. By combining these two approaches, digital twins benefit from a continuous validation pipeline that spans from early-stage software verification to hardware-level performance testing, laying the foundation for reliable predictive maintenance and operational optimization of batteries.
III Framework Design of five-tier digital twin for battery management
In this section, we present the digital twin framework for battery management applications and BMS design. First, we introduce the overall framework design. Then, we articulate several potential applications of this digital twin in predictive and prescriptive battery analysis.
Figure 1: Overview of the proposed five-tier digital twin intelligence framework, systematically integrating geometric modeling, descriptive analytics, predictive forecasting, prescriptive optimization, and autonomous control, enabling comprehensive lifecycle management and proactive decision-making for intelligent battery management.
III-A Architecture
Our five-tier digital twin framework is built upon three core technological modules that work synergistically and systematically: (i) a Virtual 3D Environment powered by NVIDIA Omniverse for geometric modeling and visualization, (ii) a multi-physics simulation engine for high-fidelity battery state estimation and prediction, and (iii) an AI engine leveraging NVIDIA PhysicsNeMo with PIML to tighten synchronization between the physical battery and its digital replica. Together, these modules enable progression across five distinct intelligence tiers, as shown in Figure 1. Each layer is constructed upon the foundation of the previous one, collectively enabling comprehensive battery lifecycle management.
From bottom to top, Tier 1 (Geometric) builds high‑fidelity 3D representations of cells, modules, and packs, implemented with a Universal Scene Description (USD)‑based environment in NVIDIA Omniverse as the spatial backbone; Tier 2 (Descriptive) binds real‑time sensor streams to this geometry to create a live, data‑rich twin that visualizes temperature distributions, voltage/current profiles, and operating conditions; Tier 3 (Predictive) leverages PIML to forecast capacity‑degradation trajectories, estimate SOH and RUL, and quantify thermal‑runaway risk under diverse scenarios; Tier 4 (Prescriptive) translates these forecasts into optimal operating strategies, such as fast‑charging protocols, cell‑balancing schedules, and cooling‑system set‑points, while enforcing safety and operational constraints; and Tier 5 (Autonomous) enables the framework to achieve closed-loop, AI-driven control, where decisions are executed independently and models adapt dynamically through continual learning. This five-tier digital twin intelligence framework design is developed based on the following three key technologies:
III-A1 Virtual 3D Environment Module
NVIDIA Omniverse serves as the foundational platform for constructing high-fidelity virtual representations of battery systems. USD files enable standardized geometric modeling spanning from individual cell components to complete battery system assemblies with many individual packs. Within this framework, this module is designed based on a USD-based 3D scene environment, which is completely programmable using a Python script to add or remove 3D objects and determine their spatial location. In our case, the scene graph captures the full battery hierarchy from cell to module to pack to represent a 20 kWh-level lithium-ion battery energy storage system. On top of this geometric structure of an energy storage system, the module adds a semantic schema that assigns meaning to each element. While the scene graph defines how cells, modules, and packs of the batteries are arranged in space, the semantic schema describes what each object represents and how it should be interpreted. For instance, a node can be marked as a temperature sensor, linked to the quantity it measures, and associated with a unit such as degrees Celsius. In this way, the schema ensures that the 3D scene is not just a geometric model but also a machine-readable map of physical roles and properties. To ensure that these semantic descriptions are usable across different scenarios in the digital twin, the module also fixes global coordinate frames, standardizes unit systems, records sensor locations, and assigns unique identifiers. These definitions make it possible to map live telemetry onto the 3D geometry without ambiguity and to exchange data consistently among different simulation and control tools.
In addition, the 3D environment module defines live data mappings that connect telemetry streams and simulation outputs to the corresponding geometric entities in the scene. These mappings specify how data is transmitted and processed, including communication topics, update rates, and data types, while providing standardized input–output interfaces to connect with other tiers. The module also exposes read/write APIs for Tier 3 predictive services to present measured, estimated, and predicted results such as SOH and RUL in the 3D scene, while for Tier 4 controllers to visualize recommended setpoints. Figure 2 illustrates this integration within the Omniverse digital twin interface. This figure presents a 3D battery pack model augmented with live sensor dashboards and overlays of predicted SOH, RUL, and temperature variations results. Through this design, the shared 3D scene serves as the central hub where all information about the operational condition, real-time system information, and optimized control commands are integrated, ensuring consistency and coordination inside the battery digital twin.
Figure 2: Battery digital twin visualization in NVIDIA Omniverse, showing pack-level monitoring with predictive analytics. The system integrates real-time health indicators with SOH, RUL, and temperature prediction modules to support proactive safety and maintenance decisions.
III-A2 Multi-physics Simulation Engine
Different from the preliminary modelling techniques that describe electrochemical, thermal, or degradation processes in isolation due to the mechanistic complexity, the simulation engine in the battery digital twin integrates these domains into a unified, co-evolving environment. It is designed to advance electrochemical, thermal, and mechanical solvers into a coupled and comprehensive model so that the digital twin reproduces battery dynamics in a physically consistent manner. For example, electrochemical reactions and ohmic resistance heat are identified as different heat generation sources that result in thermal dynamic variations. The rising temperature inside the cell then has an impact on the reaction kinetics and transport processes as described by the Arrhenius equation, reshaping voltage responses and accentuating spatial non-uniformities across cells. When mechanical effects are additionally considered, lithiation-induced swelling and assembly preload alter contact conditions and conductance pathways, thereby capturing the loss of interfacial integrity and localized resistance increases observed in real packs. By resolving these cause-and-effect loops, the multi-physics simulation engine can reproduce key phenomena such as hot-spot initiation during fast charging, thermal propagation across modules, and polarization shifts under transient loads.
This multi-physics simulation engine not only aligns with experimentally observed behaviours of actual battery systems but also produces high-quality synthetic datasets for the battery digital twin. It also conducts calibrated simulation campaigns in which operating profiles, ambient conditions, thermal management strategies, and aging states are systematically synchronised, generating multi-domain information in time series such as voltage, current, temperature fields, concentration distributions, and derived health indicators. Due to these outputs being computed in a coupled electrochemical–thermal–mechanical model, they are physically interpretable and can be used to deduce other critical latent states that are hard to observe using sensors. This enriches the battery health-related information, making it particularly valuable for supervised learning and benchmarking. Furthermore, the data reliability of this simulation engine is ensured through real-time calibration against experimental measurements to make the simulation responses match well with the observed responses. Additionally, physical consistency checks based on universal physical principles, such as energy conservation laws, are utilised to further reduce the mismatch between the digital twin battery and the physical battery.
The augmentation datasets generated from the multi-physics simulation engine support the formation and enhance the performance of digital twin by (i) providing training supplementary datasets for downstream PIML tasks, (ii) forming operational scenario libraries that reveal trade-offs among various factors, such as charging rate, round-trip efficiency, and thermal effect and (iii) verification of control policies under rare or hazardous conditions before safe transfer to HIL testbeds. By combining these techniques, the simulation engine elevates the twin from a passive mirror of sensor streams into an active experimental platform for intelligent battery management, monitoring and prognostics.
III-A3 AI Engine Module
The AI engine provides powerful support for the intelligent control and accurate state monitoring of digital-twin batteries. Compared with conventional model-based approaches, AI offers more efficient and autonomous learning capabilities, enabling knowledge transfer and improved generalization. In our digital twin framework design, NVIDIA PhysicsNeMo serves as the foundation of the AI engine, enabling the battery digital twin to deliver real-time intelligent estimation, prediction and management while preserving physical consistency and interpretability. PhysicsNeMo follows a PIML paradigm that incorporates governing principles such as charge conservation, electrochemical kinetics, and thermodynamic constraints are embedded directly into neural architectures and training objectives, ensuring both estimations and predictions remain physically plausible and trustworthy rather than purely data-driven. PhysicsNeMo implements neural PDE solvers to accelerate electrochemical and thermal dynamics by several orders of magnitude compared to classical finite-element methods. Meanwhile, transformer-based or other sequence-learning models capture long-term dependencies that characterize degradation trajectories. Together, these methods provide robust forecasts of voltage response, temperature evolution, SOH, RUL, and associated uncertainty at battery cell, module, and pack levels.
Building on its predictive accuracy, the AI engine also supports prescriptive analytics by turning predictions into concrete operating strategies. For example, reinforcement learning with experience replay utilizes past driving or charging information to gradually improve the charging strategy, and thermal management is scheduled as batteries age. Transfer learning accelerates the roll-out of new systems by reusing models trained on similar fleets, reducing the necessity of model retraining. In addition, AI-driven multi-objective optimization methods are applied to generate clear trade-off curves that show, for example, how increasing the charging rate may reduce efficiency or raise cell temperature. This information gives operators and upper-level controllers practical choices, facilitating them to select strategies that well-fit their performance and safety requirements under varying conditions.
From a system-level perspective, the AI engine drives autonomous operation by turning forecasts and optimizations into direct control actions. It utilized continual learning to adapt operational policies as usage conditions vary, and employs fault-tolerant mechanisms that keep the system functional when components fail. Before new decisions are enacted, the engine runs fast predictive simulations to conduct risk detections such as thermal runaway or internal short-circuit. To validate these learned policies work effectively on real battery systems, the AI engine is coupled with HIL platforms (e.g., OPAL-RT), which execute virtual commands against real controllers and components. In this way, optimized policies are validated under realistic constraints before deployment, ensuring that the digital twin can operate as a self-governing battery manager across its lifecycle. Overall, these capabilities establish the AI engine as the intelligence core of the digital twin, as illustrated in Figure 3.
Figure 3: Integrated Technical Architecture of the Battery Digital Twin Platform. The comprehensive technical architecture of our digital twin platform includes multi-physics modeling tools (e.g., SPM and DFN models), advanced simulation platforms, and physics-informed AI methodologies such as PhysicsNeMo. The layered structure enables systematic data integration, accurate predictive simulations, and effective prescriptive strategies, thus supporting proactive battery health management throughout its lifecycle.
III-B Potential Applications
The capabilities of our five-tier digital twin framework enable transformative applications across battery management, through two primary domains that collectively address intelligent battery operations.
III-B1 Predictive Battery Health Management
The most immediate value of the proposed battery digital twin lies in its ability to transform health prediction into a continuous and adaptive process. Instead of relying on sparse snapshots, SOH and RUL are evaluated in real time, adapting to variations in usage and environment. This approach improves predictive accuracy and produces risk-aware outputs expressed as probability distributions rather than deterministic forecasts, enabling operators to plan maintenance proactively with quantified confidence. Such predictive intelligence ensures that maintenance actions are both timely and cost-effective, reducing unexpected downtime and improving overall system reliability.
Beyond prediction, the digital twin enables early detection of degradation far in advance of conventional monitoring tools. By identifying subtle indicators of failure, such as cell imbalance or incipient thermal runaway, it allows operators to isolate or rebalance systems before issues escalate. This capability supports practical applications across fleets and grid storage by preventing propagation of failures and guiding warranty or replacement decisions based on emerging degradation patterns. By distinguishing short-term performance drift from irreversible damage, the system ensures that minor fluctuations are managed efficiently while critical risks are addressed without delay, strengthening both safety and operational resilience.
III-B2 Battery Repurposing and Value Maximization
Beyond health prediction, the digital twin enables actionable strategies to maximize battery value across its entire lifecycle. During operation, forecasts are translated into adaptive charging protocols that shorten charging duration while keeping temperature rise and energy loss within safe limits, and intelligent thermal management that dynamically adjusts airflow and coolant flow according to predicted heat loads. These strategies not only improve energy efficiency but also mitigate thermal stress that accelerates degradation. At the system level, the twin orchestrates load distribution across modules with different health states. Instead of uniform current sharing, it redistributes demand away from overstressed or degraded packs, thereby protecting weaker components while extracting maximum capacity from healthier ones. This capability is particularly valuable in grid-scale storage plants integrating batteries of mixed chemistries and vintages, where conventional rule-based control often leads to underutilization of robust modules and premature aging of already stressed ones [12].
The framework further supports autonomous lifecycle management by combining health forecasts with economic reasoning. Instead of fixed service intervals, it continuously evaluates the cost–benefit trade-offs of interventions such as rebalancing or component replacement, ensuring maintenance actions are both timely and economically justified. Finally, by preserving complete degradation histories, the digital twin provides accurate assessments of residual capacity at end-of-life. This enables reliable decisions for second-life deployment in stationary storage or recycling, ensuring safe, efficient, and sustainable reuse while maximizing residual value.
IV Case Study: Integrated Multi-physics Calibration and Uncertainty-quantified SOH Prediction
To validate the effectiveness of the proposed framework, we present a case study on multi-physics calibration of cell voltage and temperature, and uncertainty-quantified SOH prediction for battery operation.
IV-A Problem Settings
This case study demonstrates the practical application and effectiveness of our proposed digital twin framework for intelligent battery management using the XJTU battery dataset [13]. The dataset consists of run-to-end of life experiments performed on 55 cylindrical 18650 Nickel-Cobalt-Manganese (NCM) lithium-ion batteries ($\mathrm{LiNi_{0.5}Co_{0.2}Mn_{0.3}O_{2}}$), each with a nominal capacity of $2000\,\mathrm{mAh}$ and nominal voltage of $3.6\,\mathrm{V}$. The experiments are carried out using six different charge-discharge regimes designed to emulate various realistic operational conditions, ranging from constant C-rate cycling and variable discharge conditions to randomized and specialized satellite battery usage patterns.
In our study, we combine physical modeling with data-driven models to address two core functions of practical battery management. First of all, we employ a DFN electrochemical model combined with a lumped thermal model to simulate the battery operation. Accurate calibration of this multi-physics model is the key to reliably simulating the voltage and temperature behavior of batteries under diverse cycling conditions, as represented in the dataset. Calibrating these models is fundamental, as precise physical simulations constitute the cornerstone of an effective digital twin for advanced battery management, as we addressed before. Second, utilizing the calibrated simulation data sets up the foundation to develop an advanced predictive model to accurately predict SOH while simultaneously quantifying the uncertainty. The integration of uncertainty quantification addresses a critical gap in existing deterministic prediction methodologies, enabling risk-informed operational decision-making for improved battery lifecycle management.
IV-B Model Architecture
Our integrated battery digital twin combines electrochemical-thermal simulations, Physics-informed Neural Networks (PINNs), and uncertainty quantification through Deep Autoencoding Gaussian Mixture Models (DAGMM). This approach enables accurate and physically consistent predictions of battery SOH and robust quantification of prediction uncertainty.
IV-B1 Multi-physics Modeling and Calibration
We integrate the SPM and DFN electrochemical models with a lumped thermal model using PyBaMM as the model source platform [14]. The SPM offers computationally efficient simulations suitable for rapid, real-time predictions, while the DFN model provides detailed modeling of internal lithium-ion transport processes and reaction kinetics. Lumped thermal model is incorporated with the DFN model to ensure an accurate representation of thermal dynamics that effectively captures dynamic battery behavior under various operational conditions.
To ensure the fidelity of our simulation models with respect to actual battery performance, we further implement a rigorous Bayesian optimization procedure for parameter calibration. This optimization systematically adjusts over 15 model parameters, such as diffusivities, reaction rate, conductivities, electrode geometries, and thermal conductivities, by minimizing discrepancies between simulated outputs and experimentally measured voltage and temperature data. Through iterative refinement using Bayesian optimization methods, our calibrated multi-physics models attain strong consistency with observed battery responses, ensuring reliable simulation outcomes.
IV-B2 Physics-informed Neural Network
Leveraging NVIDIA PhysicsNeMo, we construct PINNs that combine deep learning approaches with PyBaMM simulations. These PINNs are trained using data from our calibrated multi-physics simulations, incorporating a physics-informed loss function as the learning bias that integrates both data-driven accuracy terms and physical constraints. Specifically, the loss function penalizes deviations from fundamental battery physics and thermal equilibrium. By enforcing these constraints via automatic differentiation at strategically selected points within the battery simulation domain, this proposed PINN approach effectively captures the complex battery dynamics while maintaining adherence to underlying physical laws.
IV-B3 Uncertainty Quantification via DAGMM
In addition to these approaches mentioned above, our predictive model further incorporates DAGMM for uncertainty quantification. DAGMM leverages an autoencoder network to transform operational data into latent representations, subsequently modeled by a Gaussian Mixture Model (GMM). The uncertainty metric is obtained through an energy score, evaluating latent representations’ likelihood within the learned GMM. This energy-based score identifies distributional anomalies indicative of prediction uncertainty.
IV-C Evaluation Results
Rigorous evaluations are given to demonstrate the superior predictive accuracy and reliability of the proposed digital twin framework through multi-physics model calibration and uncertainty quantifications.
Figure 4: Multi-physics calibration results against the XJTU battery dataset after Bayesian optimization. (a) Voltage prediction under 1C, 2C, and 3C discharge rates. (b) Temperature prediction under the same operating conditions.
For the multi-physics calibration, the electrochemical-thermal model demonstrated good estimation accuracy compared with the XJTU battery dataset with the use of Bayesian optimization as shown in Figure 4. For voltage prediction, the calibrated model achieved an MAPE of 0.92% (1C), 1.06% (2C), and 1.57% (3C). For temperature prediction, the MAPE reached 0.07% (1C), 0.18% (2C), and 0.39% (3C). These results validate the effectiveness of the multi-physics model that can accurately capture both potential and thermal responses of batteries under different operational regimes.
Figure 5: Prediction results across six charging protocol datasets. Each panel plots battery SOH versus cycle index for a distinct dataset (2C, 3C, R2.5, R3, RW, Satellite). The red curve shows the true SOH labels and the blue curve shows predictions from our model. Panels share comparable axis limits to enable fair visual comparison across datasets. The results indicate that predictions closely follow the true trajectories under both constant rate and randomized profiles.
Building upon the calibrated simulations, we can further assess the predictive performance of our PINN-based SOH prediction model, including its capability to quantify prediction uncertainty. As shown in Figure 5, our model demonstrated robust predictive accuracy in SOH, achieving mean absolute percentage errors (MAPE) consistently below 3%. Most importantly, the integrated uncertainty quantification via the DAGMM provided reliable indications of prediction confidence. As shown in Figure 6, the energy-based uncertainty scores exhibited a strong positive correlation with the actual prediction errors, effectively highlighting scenarios where predictions were less reliable due to data distribution shifts.
Figure 6: Impact of noise injection on prediction error and the proxy metric for uncertainty.
These comprehensive evaluation outcomes validate that our integrated digital twin framework not only achieves precise physical modeling and accurate health predictions but also effectively quantifies uncertainties, significantly enhancing the robustness and reliability of battery management decisions.
V Future Directions
Based on the initial results from the case study mentioned above, we discuss several future directions that can be explored to further enhance the intelligent battery management using this digital twin architecture.
Foundation Models for Battery Intelligence: Large Language Models (LLMs) are foundation models trained on massive text corpora using transformer architectures, enabling diverse tasks such as reasoning, summarization, and question answering without task-specific supervision. Beyond language, they serve as general-purpose engines for knowledge representation and code generation across scientific domains. Building on these capabilities, LLMs adapted on battery-related literature present promising opportunities for digital twins through automated knowledge extraction and model generation [15]. Future implementations envision specialized battery foundation models that automatically design PINN architectures for specific chemistries, synthesize insights from vast research to identify degradation mechanisms, recommend experimental protocols, and support conversational interfaces for intuitive querying of battery states and explanations of complex degradation phenomena.
Blockchain-based Battery Passport Systems: The implementation of blockchain in battery lifecycle management offers transformative potential for future battery passport systems, enabling complete traceability and transparency. Distributed ledger technologies will record manufacturing data, operational history, maintenance, and performance metrics as immutable entries, ensuring reliable information sharing across manufacturers, operators, and recyclers. Smart contracts could automate management decisions such as maintenance scheduling and end-of-life processing, reducing fraud in condition reporting. Blockchain-based passports will further support circular economy initiatives by providing verified health data for second-life applications, allowing batteries retired from vehicles to be reused in stationary storage with confidence.
Differentiable Simulation for Optimal Control: The development of fully differentiable battery simulation frameworks enables gradient-based optimization of battery operational strategies directly through physics-based models, eliminating the need for computationally expensive reinforcement learning approaches. Future implementations could leverage automatic differentiation through simulation platforms to enable direct optimization of charging protocols, thermal management strategies, and load balancing decisions with respect to battery health and performance objectives. Differentiable simulation supports real-time model predictive control that optimizes battery operation while accounting for complex multi-physics interactions and degradation constraints. This approach provides enhanced transparency compared to black-box optimization methods while enabling principled handling of operational constraints through physics-based penalty terms in the optimization objective.
VI Conclusion
In this article, we presented an intelligent battery management concept based on a five-tier digital twin framework for autonomous operations. We first highlighted current BMS challenges and the need for advanced digital twin approaches. Then, we introduced the proposed framework, progressing from geometric modeling to autonomous operation, supported by PIML and uncertainty quantification. Its applications span predictive health management, operational optimization, and lifecycle control. Finally, a case study validated our design, showing sub-1% voltage/temperature errors and robust SOH predictions with MAPE below 3%, demonstrating the framework’s potential for interpretable modeling and reliable decision-making.